If an investor had a $25,000 long-term capital gain on a $100,000 investment from 1984 to 2010, her real rate of return was most likely

a.

zero.

b.

negative.

c.

equal to the expected rate of inflation.

d.

equal to the nominal rate of inflation.
Answer #1

Answer

Option b

It is negative as inflation is higher than the nominal rate in this period in the world

nominal return is

i=(FV/PV)^(1/n)

n=years =2010-1984=26

i=(125000/100000)^(1/26)-1

=0.00861938

=0.86%

the normal inflation is above 2% in this time so it is negative